Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2017-09-28

Observers: Kris Rybicki, Simon Hodgkin
TAT: Regis Cartier, Janet Chen
Support team: Morgan Fraser, Emma Callis

Calibrations: biases, dome arcs+flats - apparently some problem with flats

Conditions: Quite cloudy - but opened just after sunset.

00:30 Seeing ~1 arcsec
03:10 Clouds! Carrying on...
05:29 Clouds! Closed
07:20 Reopened (still quite grotty)


UT       Target            Type            Comment
==============================================================================================     
00:00    AT2017gxs         ACQ             Nothing in image
00:15    ATLAS17lai        ACQ             Nothing in image
00:27    EG131             STD
00:56    AT2017gah         GR11+GR16       Follow-up
02:15    AT2017gpp         GR13            Follow-up
03:12    Feige110          STD
03:33    Gaia17cib         GR13            Class: SNIa at +26d
03:57    ATLAS17kwg        GR13            Class: SNIa
04:16    DES17C1ald        GR13            Class: SNIa
04:48    AT2017guh        GR11+GR16      Follow-up
07:20    AT2017gci         GR13            Follow-up
08:08    AT2017jbu         GR11+GR16       Follow-up
09:??    GD71              STD
Comments